2) Two charges, each with 15.0 uC but with opposite sign are 33.0 mm apart. What is the magnitude of the dipole moment (magnitude) in SI units.

A)2.07*10^-7

B)6.82*10^-9

C)4.95*10^-7

D)7.43*10^-12

E)9.90*10^-7

The correct answer us C but could you please show work. Thanks!
Answer #1

Dipole moment= charge on either end × distance between two ends

Dipole moment = 15 × 10 -6  × 33 × 10 -3

Dipole moment = 495 × 10 -9

Dipole moment = 4.95 × 10 -7

Correct option is C